Please do it clear so I can do it well.
4. value 10.00 points The Pet Store experienced the following events for the 2016 accounting perlocd 1. Acquired $60,000 cash from the issue of common stock 2. Purchased $65,000 of in on account 3. Received goods purchased in Event 2 FOB shipping point; freight cost of $900 paid in cash. 4. Sold inventory on account that cost $38,000 for $71,000 Event 4 was $620. The goods were shipped FOB destination. was paid for the freight cost. 6. Customer in Event 4 returned $4,200 worth of goods that had a cost of $2,150. 8. Paid $59,200 cash on 9. Paid $2,600 for advertising expense. 10. Paid $3,100 cash for insurance expense. $58,300 cash from accounts receivable. payable Required a. Which of these events affect period (selling and administrative) costs? Which result in product costs? If neither, label the transaction NA Transaction Cost 10.
in the for Expense Net Flow Cash+ 60,000 3. Freight 10 Pd
Answer #1

Ans-a-

Transaction Cost
1 NA
2 NA
3 Product
4 Product
5 Period
6 Product
7 NA
8 NA
9 Period
10 Period

  

THE PET STORE

Horizontal Statement Model for 2016

Balance Sheet Income Statement
Assets ($) = Liabilities ($) + Stockholder's Equity ($) Revenue ($) - Expenses ($) = Net Income ($) Statement of Cash Flows ($)
Cash + Accounts Receivable + Inventory = Accounts Payable + Common Stock + Retained Earnings
1- Issue of common stock 60,000 + + = + 60,000 + - = 60,000 FA.
2- Purchase of Inventory + + 65,000 = 65,000 + + - =
3-Freight -900 + + 900 = + + - = (900) OA
4-a--Sale of inventory + 71,000 + = + + 71,000 71,000 - = 71,000
4-b- Cost of goods sold + + -38,000 = + + -38,000 - 38,000 = -38,000
5- Freight cost -620 + + = + + -620 - 620 = -620 (620) OA
6-a- Sales Return + -4,200 + = + + -4,200 -4,200 - = -4,200
6-b- Return Inventory + + 2,150 = + + 2,150 - -2,150 = 2,150
7- Collection on accounts receivable   58,300 + -58,300 + = + + - = 58,300 OA
8- Payment on account -59,200 + + = -59,200 + + - = (59,200) OA
9- Payment of advertising -2,600 + + = + + -2,600 - 2,600 = -2,600 (2,600)OA
10-Payment of insurance -3,100 + + = + + -3,100 - 3,100 = -3,100 (3,100) OA
Ending Balance $51,880 + $8,500 + $30,050 = $5,800 + $60,000 + $24,630 $66,800 - $42,170 = $24,630 $51,880 NC

OA- Operating Activities

FA- Financial Activities

NC-Net change in cash